plane travels 175 miles with the wind in the same amount of time it flew 150 miles against the wind. what was the speed of the plane in the still air if the wind was blowing at 20 mph?
Answer by mananth(16949)

plane travels 175 miles with the wind in the same amount of time it flew 150 miles against the wind. what was the speed of the plane in the still air if the wind was blowing at 20 mph?
..
let speed of plane in still air be x mph
wind speed = 20 mph
..
speed with wind = x+20: d=175 miles
against wind speed = x-20: d= 150 miles
...
175/(x+20)=150(x-20)
175(x-20)=150(x+20)
175x-3500=150x+3000
reorganise
25x=6500
/25
x=6500/25
x=260 mph speed of plane in still air
...
m.ananth@hotmail.ca